OpenStudy (theedwardsfamily):

Help Please!!!!!!!!!!!!!!!! Points C and D lie on graph of the equation y = 1.3x. What are the y-coordinates for each point? Point C: (3, ___) Point D: (−6, ___)

OpenStudy (campbell_st):

so the equation is 1. y = 1.3x or 2. \[y = 1.3^x\]

OpenStudy (theedwardsfamily):

number 1

OpenStudy (theedwardsfamily):

@campbell_st

OpenStudy (campbell_st):

ok... so just replace x with 3 to find y then x with -6 remember the equation can be written as \[y = 1.3 \times x\] so for x = 3 its \[y = 1.3 \times 3\] I'll leave you to calculate the answers

OpenStudy (theedwardsfamily):

3.9 @campbell_st ?

OpenStudy (campbell_st):

that's correct so you have the point (3, 3.9) now you need to do the same, replacing x with - 6

OpenStudy (theedwardsfamily):

ok i get it.